2.21 Among donors at a blood center, 1 in 2 gave type O+ blood, 1 in 11 gave O~, 1 in 4 gave A and 1 in 20 gave AT. What is the probability that the first person who shows up tomorrow to donate blood has the following blood type? a Type O b Type O C Type A d Either type AT or O e Neither type A+ nor O+
Answer #1

(a) Prob. = 1/2 = 0.50
(b) Prob. = (1 + 1)/(2 + 11) = 0.1538
(c) Prob. = (1 + 1)/(20 + 4) = 0.0833
(d) Prob. (A- \cup O-) = Prob.(A-) + Prob.(O-) = (1/20) + (1/11) =  0.1409
(e) Prob. ((A+)c  \cap (O+)) = 1 - Prob.(A+) - Prob.(O+) = 1 - (1/4) - (1/2) = 0.25